a) The number of grams of a certain radioactive substance present after t hours is given by the equation Q = Q0e−0.45t, where Q0 represent the ini- tial number of grams. How long will it take 2500 grams to be reduced to 1250.


1
Expert's answer
2021-12-15T10:06:03-0500

"\displaystyle
\text{Given }\\
Q(t)=Q(0)e^{-0.45t}\\
\text{We solve for t to obtain}\\
e^{-0.45t}=0.5\\
\text{Taking the log of both sides, we have}\\
-0.45t=\ln (0.5)=-0.6931\\
\implies t = 1.54 hrs"